OpenStudy (anonymous):

Help! Some Polynomials questions? What is the correct simplification of the expression (3^a/b^7)^2? What is the correct simplification of the expression 6b/a^-9? What is the correct simplification of the expression -5x^2(4x - 6x^2 - 3)?

OpenStudy (anonymous):

please i really need help!

OpenStudy (anonymous):

\[({3^a \over b^7})^2={(3^a)^2\over (b^7)^2}={3^{2a} \over b^{14}}\] \[{6b \over a^{-9}}=6b \cdot a^9 = a^96b\] \[-5x^2(4x - 6x^2 - 3)=-20x^3+30x^4+15x^2\] I don't know if the last one can be considered a semplification, but I think the others are ok